# The stellar bar – dark matter halo connection in the TNG50 simulations

### v2

@@ Appendix C Estimation of $R_{\star,60}$, $R_{\star,90}$ and $R_{200}$
Most of the disks in our sample are well within the sphere of 30 kpc radius. Hence it is reasonable to estimate the radius containing 30%, 60% and 90% of the stellar mass within a sphere of radius 30 kpc. Some of the disks in our sample have small values of $R_{\star,30}$ and $R_{\star,60}$, and estimation the DM halo properties at small radii can suffer from issues of numerical relaxation (Power et al., 2003). Following the method presented in Power et al. (2003), we estimate the radius that bounds the volume inside which the two-body relaxation time scale $t_{relax}$ is smaller than the age of the Universe $t_{age}$. The radius at which $t_{relax}=t_{age}$, can be adopted as the radius of convergence ($r_{conv}$) of the inner region of the DM halo. $r_{conv}$ depends on the number of DM particles ($N(<r)$), the mean density ($\bar{\rho}(r)$) and the critical density of the Universe at redshift $z_{r}$ ($\rho_{crit}(z_{r})=3H^{2}(z_{r})/8\pi G$), and is a solution to the following equation:
@@ Appendix D Median halo spin for galaxies after relaxing the fixed DM mass criterion
We calculate the halo spin at multiple radii of the DM halo, namely $r_{m}=5,$ and 10 kpc in Section 2.1 and 5. We also determine the halo spin at $R_{\star,60}$ and $R_{\star,90}$, the radius at which the stellar mass reaches 60% and 90% of the total mass inside a spherical region (centered at disk center) of radius 30 kpc. To estimate $R_{\star,X}$, we choose a maximum radius of 30 kpc as all of the galaxies in our sample are well within the 30 kpc radius. In addition, we estimate the halo spin at the virial radius $R_{200}$, far from the central disk and close to the disk at $0.15R_{200}$ and $0.05R_{200}$, where we expect a greater impact of the disk on the surrounding DM particles. We present the detailed method of estimation of each radius in Appendix C.
Figure 14: The connection between bar strength and dark matter halo spin at redshift $z_{r}=$0.0 (top row), 0.1 (middle row) and 1.0 (bottom row). Each panel shows the median halo spin and median bar strength (circles, squares and $\star$”) for each of the bar strength groups from Table 1, for different radii. In all panels, green circles are for the smallest radii, magenta squares are for the next radii and blue stars are for the largest radii, with the shaded regions of similar colors bounded by the 16th and 84th percentile curves of halo spin distribution. Left column shows halo spin at $r=5$, 10 and 100 kpc; middle column for $r=0.05R_{200},0.15R_{200}$ and $0.5R_{200}$; the right column for $r=R_{\star,60}$ and $R_{\star,90}$ (see Section 6.1). At low redshift and small radii, the median halo spin decreases with increasing bar strength. The shape of the shaded regions highlights the skewed nature of the spin distribution in each bar strength group.
@@
The halo spin at higher redshifts (Figure 14, middle row and bottom row), for example at $z_{r}=1.0$ is very different from halo spin at $z_{r}=0.0$, although the spin distribution at $z_{r}=0.1$ has more similarity with spin at $z_{r}=0.0$. There is significantly less number of low-mass galaxies that settle into a disk in our high redshift samples and unlike $z_{r}=0.0$, we observe a higher median halo spin of the unbarred galaxies at $z_{r}>0.0$. Overall, the halo spins for unbarred galaxies ($0.20<A_{2}/A_{0}<$ 0.2) are higher compared to the barred galaxies ($0.70.2<A_{2}/A_{0}<$ 0.7) even at $z_{r}=0.1$. However, the low redshift trend does not hold at high redshift. At $z_{r}=1.0$, the median halo spin close to the disk (green and magenta lines) is either similar for barred and unbarred galaxies or the trend reverses at very low radii (green line in panel g; magenta line in panel i). At large radii, away from the disk (blue lines in panels g and h), the halo spin is higher for unbarred galaxies compared to barred galaxies.
@@ Appendix E Mass fraction in the disk within $R_{\star,90}$
Figure 15: Stellar mass fraction ($M_{\star}/M_{tot}$; left column), dark matter mass fraction ($M_{DM}/M_{tot}$; middle column) and gas mass fraction ($M_{gas}/M_{tot}$; right column) at redshift $z_{r}=$0.0 (top row), 0.1 (middle row) and 1.0 (bottom row), within spherical radius $R_{\star,90}$ and $|z|<2$ kpc. In each panel, reddish/blue/green points are values of each quantity in different bar strength groups (Table 1). Blue lines connect the median values (dark blue circles) in each bar strength group and the shaded regions are bounded by the 16th and 84th percentile of each of the above quantities in each bar strength group.
Figure 15 shows the stellar mass fraction, the DM mass fraction and the gas mass fraction within radius $r<R_{\star,90}$ and $|z|<2$ kpc from the disk mid-plane, of all our sample galaxies at different redshifts. In Figure 15, we present the mass fractions of all galaxies at redshift $z_{r}=0.0$ (top row), 0.1 (middle row) and 1.0 (bottom row). The left column shows the stellar mass fraction ($M_{\star}/M_{tot}$; orange circles), middle column is for DM mass fraction ($M_{DM}/M_{tot}$; blue circles) and the right column is for the gas mass fraction ($M_{gas}/M_{tot}$; green circles), where $M_{tot}$ is the combined mass of all components. The dark-blue big solid circles show the median mass fraction in each panel and the shaded region is bounded by the 16th percentile and the 84th percentile curves.

### v3

@@ Appendix C Estimation of $R_{\star,60}$, $R_{\star,90}$ and $R_{200}$
Most of the disks in our sample are well within the sphere of 30 kpc radius. Hence it is reasonable to estimate the radius containing 30%, 60% and 90% of the stellar mass within a sphere of radius 30 kpc. Some of the disks in our sample have small values of $R_{\star,30}$ and $R_{\star,60}$, and estimation the DM halo properties at small radii can suffer from issues of numerical relaxation. Following the method presented in Power et al. (2003), we estimate the radius that bounds the volume inside which the two-body relaxation time scale $t_{relax}$ is smaller than the age of the Universe $t_{age}$. The radius at which $t_{relax}=t_{age}$, can be adopted as the radius of convergence ($r_{conv}$) of the inner region of the DM halo. $r_{conv}$ depends on the number of DM particles ($N(<r)$), the mean density ($\bar{\rho}(r)$) and the critical density of the Universe at redshift $z_{r}$ ($\rho_{crit}(z_{r})=3H^{2}(z_{r})/8\pi G$), and is a solution to the following equation:
@@ Appendix D Median halo spin for galaxies after relaxing the fixed DM mass criterion
We calculate the halo spin at multiple radii of the DM halo, namely $r_{m}=5,$ and 10 kpc in Section 3 and 5. We also determine the halo spin at $R_{\star,60}$ and $R_{\star,90}$, the radius at which the stellar mass reaches 60% and 90% of the total mass inside a spherical region (centered at disk center) of radius 30 kpc. To estimate $R_{\star,X}$, we choose a maximum radius of 30 kpc as all of the galaxies in our sample are well within the 30 kpc radius. In addition, we estimate the halo spin at the virial radius $R_{200}$, far from the central disk and close to the disk at $0.15R_{200}$ and $0.05R_{200}$, where we expect a greater impact of the disk on the surrounding DM particles. We present the detailed method of estimation of each radius in Appendix C.
Figure 14: The connection between bar strength and dark matter halo spin at redshift $z_{r}=$0.0 (top row), 0.1 (middle row) and 1.0 (bottom row). Each panel shows the median halo spin and median bar strength (circles, squares and $\star$”) for each of the bar strength bins from Table 1, for different radii. In all panels, green circles are for the smallest radii, magenta squares are for the next radii and blue stars are for the largest radii, with the shaded regions of similar colors bounded by the 16th and 84th percentile curves of halo spin distribution. Left column shows halo spin at $r=5$, 10 and 100 kpc; middle column for $r=0.05R_{200},0.15R_{200}$ and $0.5R_{200}$; the right column for $r=R_{\star,60}$ and $R_{\star,90}$ (see Section 6.1). At low redshift and small radii, the median halo spin decreases with increasing bar strength. The shape of the shaded regions highlights the skewed nature of the spin distribution in each bar strength bin.
@@
The halo spin at higher redshifts (Figure 14, middle row and bottom row), for example at $z_{r}=1.0$ is very different from halo spin at $z_{r}=0.0$, although the spin distribution at $z_{r}=0.1$ has more similarity with spin at $z_{r}=0.0$. There is significantly less number of low-mass galaxies that settle into a disk in our high redshift samples and unlike $z_{r}=0.0$, we observe a higher median halo spin of the unbarred galaxies at $z_{r}>0.0$. Overall, the halo spins for unbarred galaxies ($0.20<A_{2}/A_{0}<$ 0.2) are higher compared to the barred galaxies ($0.70.2<A_{2}/A_{0}<$ 0.7) even at $z_{r}=0.1$. However, the low redshift trend does not hold at high redshift. At $z_{r}=1.0$, the median halo spin close to the disk (green lines in panel e, and magenta and blue lines in panel f) is either similar for barred and unbarred galaxies or the trend reverses at very low radii (magenta line in panel f). At large radii, away from the disk (blue and magenta lines in panel e), the halo spin is higher for unbarred galaxies compared to barred galaxies.
@@ Appendix E Mass fraction in the disk within $R_{\star,90}$
Figure 15: Stellar mass fraction ($M_{\star}/M_{tot}$; left column), dark matter mass fraction ($M_{DM}/M_{tot}$; middle column) and gas mass fraction ($M_{gas}/M_{tot}$; right column) at redshift $z_{r}=$0.0 (top row), 0.1 (middle row) and 1.0 (bottom row), within spherical radius $R_{\star,90}$ and $|z|<2$ kpc. In each panel, reddish/blue/green points are values of each quantity in different bar strength bins (Table 1). Blue lines connect the median values (dark blue circles) in each bar strength bin and the shaded regions are bounded by the 16th and 84th percentile of each of the above quantities in each bar strength bin.
Figure 15 shows the stellar mass fraction ($M_{\star}/M_{tot}$; orange circles in left column), the DM mass fraction ($M_{DM}/M_{tot}$; blue circles in middle column) and the gas mass fraction ($M_{gas}/M_{tot}$; green circles in right column) within radius $r<R_{\star,90}$ and $|z|<2$ kpc from the disk mid-plane, of all our sample galaxies at redshifts $z_{r}=0.0$ (top row), 0.1 (middle row) and 1.0 (bottom row). Here $M_{tot}$ is the combined mass of all components. The dark-blue big solid circles show the median mass fraction in each panel, and the shaded region is bounded by the 16th percentile and the 84th percentile curves.